OpenStudy (anonymous):

HELP!!! Will fan and medal! Solve 8 < 2b OR 2b + 15 < 5 a. b > 4 OR b < 5 b. b < 2 OR b > 3 c. b > 1 OR b < 3 d. b > 3 OR b < 4

OpenStudy (anonymous):

Divde the first one by 2. S it becomes 8/2=4 ....4<b. or. Subtract by 15 so its 5-15=10 and then divode ny 2.......10/2=5 ....b<5 .... answer is A

OpenStudy (anonymous):

thank you!